# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

### v1

@@ 5.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
Therefore, all probability measures $Q^{\prime}\in\mathcal{Q}$ that satisfy the above bound are potential minimisers of $T_{n}(\cdot)$. This is precisely the set $\mathcal{R}^{\star}_{n}$. Furthermore, by Assumption 4 we know that $\forall Q^{\prime}\in\mathcal{Q}$ we have $\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]\geq C$, so the LHS is non–negative, and since $P_{n}^{\star}$ is in this set, it is non–empty.
For the moreover part, we additionally suppose that $\mathcal{Q}$ is weak closed and that the loss is lower semi–continuous and coercive. It immediately follows that the functional
$J_{L_{n}}(Q):=\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]$
@@
is lower semi–continuous, note that this follows by the functional being over probability measures and the loss being lower bounded and lower semi–continuous [11, 4]. We then know by Lemma 2 that the sublevel sets of the functional $J_{L_{n}}(\cdot)$ are tight. Hence, by the lower semi–continuity of the functional, $\mathcal{Q}$ being weak closed, and Prokhorov’s theorem [8], the set $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq k\}$ is weak compact.
@@
Therefore, we may apply this to $\mathcal{R}_{n}^{\star}$ by picking $k=M(n\beta)^{-1}+\inf_{\nu\in\mathcal{Q}}J_{L_{n}}(\nu)$, which exists and is finite by Assumption 4, establishing that $\mathcal{R}_{n}^{\star}$ is weak compact as desired. ∎
5.2 Existence and Uniqueness of a Generalised Variational Inference Posterior
We first prove existence of GVI minimisers and then show that if a convex combination is in $\mathcal{Q}$ then it will also be in $\mathcal{R}_{n}^{\star}$ from which uniqueness then immediately follows.
Proof of Theorem 2.
By Theorem 1 we have that $Q_{n}\in\mathcal{R}_{n}^{\star}$ which is weak compact. Then, as the divergence is lower semi–continuous, we have that the functional
$T_{n}(\nu):=nJ_{L_{n}}(\nu)+\frac{1}{\beta}D(\nu:\Pi)$
is lower semi–continuous [11]. Then, as the loss is lower bounded, and the divergence is non–negative, it follows that $T^{\star}:=\inf_{\nu\in\mathcal{Q}}T_{n}(\nu)>$-$\infty$, and we may construct an arbitrary sequence $\ell_{k}:=T_{n}(\nu_{k})$, where $(\nu_{k})_{k}\subset\mathcal{R}_{n}^{\star}$, which satisfies $\ell_{k}\rightarrow T^{\star}$ as $k\rightarrow\infty$.
As $\mathcal{R}_{n}^{\star}$ is weak compact, and the functional $\nu\mapsto T_{n}(\nu)$ is lower semi–continuous, it follows that the sequence $(\nu_{k})_{k}$ is tight. Therefore, by Prokhorov’s theorem, after passing to a subsequence (without relabelling) we find that $\nu_{k}\overset{w^{\star}}{\rightarrow}\nu_{\star}$ for some $\nu_{\star}\in\mathcal{R}_{n}^{\star}$ and for which
$\liminf_{k}\;T_{n}(\nu_{k})\geq T_{n}(\nu_{\star})\,.$
Further note that $\liminf_{k}\,T_{n}(\nu_{k})=T^{\star}$ and as $T^{\star}$ is a global minimum of the GVI objective, it follows that $T_{n}(\nu_{\star})=T^{\star}$, so $\nu_{\star}\in\mathcal{Q}$ exists which minimises Equation 3. ∎
Now, we provide the proof of Corollary 3 from which Corollary 4 is an immediate consequence, as the GVI functional under the assumptions of Corollary 4 is strictly convex, and hence a minimiser needs to be a unique extreme point of $\mathcal{R}_{n}^{\star}$.
Proof of Corollary 3.
The forward direction is immediate, given that $\mathcal{R}^{\star}_{n}$ is a subset of $\mathcal{Q}$.
For the reverse direction, assume there exists $Q,P\in\mathcal{R}^{\star}_{n}$, and assume that for some $a\in(0,1)$ it holds that $aQ+(1-a)P\in\mathcal{Q}$. Then we have that, as $J_{L_{n}}(\cdot)$ is a linear functional,
$J_{L_{n}}(aQ$ $+(1-a)P)-J_{L_{n}}(P_{n}^{\star})$
$=$ $aJ_{L_{n}}(Q)+(1-a)J_{L_{n}}(P)-J_{L_{n}}(P_{n}^{\star})$
$=$ $a\left[\underbrace{J_{L_{n}}(Q)-J_{L_{n}}(P_{n}^{\star})}_{\leq M(n\beta)^{-1}}\right]+(1-a)\left[\underbrace{J_{L_{n}}(P)-J_{L_{n}}(P_{n}^{\star})}_{\leq M(n\beta)^{-1}}\right]$
$\leq$ $M(n\beta)^{-1}$
Hence, $aQ+(1-a)P\in\mathcal{R}^{\star}_{n}$ and the reverse implication holds. ∎
5.3 Support of the Prior is not required
Proof of Proposition 5.
Suppose that $\exists Q\in\mathcal{Q}$ such that both:
  1. 1.
    $Q$ and $\Pi$ are mutually singular, and
  2. 2.
    $T_{n}(Q)\leq nC+\beta^{-1}M$, for $C$ as in the proof of Theorem 1,
hold. Then, as by Assumption 3, and the definition of a divergence $0\leq D(Q:\Pi)\leq M$. Either $D(Q:\Pi)=M$, in which case $Q$ minimises the loss functional $J_{L_{n}}(\cdot)$, or $D(Q:\Pi)=k<M$, in which case $Q$ is still an element of $\mathcal{R}^{\star}_{n}$. Moreover, we have that
$nJ_{L_{n}}(Q)+\beta^{-1}D(Q:\Pi)=nJ_{L_{n}}(Q)+\beta^{-1}k\leq nC+\beta^{-1}M$
$\implies n\Bigl[J_{L_{n}}(Q)-J_{L_{n}}(P_{n}^{\star})\Bigr]\leq\beta^{-1}(M-k)$
as desired.∎
5.4 Asymptotics and Frequentist Consistency
In order to prove this result we need to collect some auxiliary results that will establish the final result.
5.4.1 Charactering the loss minimiser
We begin with a general result for $\mathcal{Q}=\mathcal{P}(\Theta)$.
@@ 5.4.1 Charactering the loss minimiser
Suppose that $f:\Theta\rightarrow\mathbb{R}$ is a $\mathcal{T}$–measurable, lower–bounded, lower semi–continuous, and coercive function. Any minimiser of the functional $J_{f}(\cdot)$ must put full measure on any measurable set containing the minimisers of the function $f(\theta)$.
@@
As the function is lower bounded, lower semi–continuous and coercive, it follows that there exists some minimiser away from the boundary of $\Theta$, specifically it achieves its minimum on some compact subset of $\Theta$ away from the boundary, so $N_{0}:=\{\theta\in\Theta:f(\theta)=f(\Theta)\}$ is well defined and non–empty. Note that this is measurable as we may define it as a countable intersection of measurable sets.
Then for any measure $\nu\in\mathcal{P}(\Theta)$ for which $\nu N_{0}=1$ the lower bound above is achieved. Hence, $\forall\varepsilon>0$ and $\forall\mu\in\mathcal{P}(\Theta)$ for which $\mu N_{\varepsilon}<1$ holds, it follows that $J_{f}(\nu)<J_{f}(\mu)$. We may see this as
$J_{f}(\mu)$ $=\int_{N_{\varepsilon}}fd\mu+\int_{N_{\varepsilon}^{c}}fd\mu$
$=\int_{\{f\leq f(\Theta)+\varepsilon\}}fd\mu+\int_{\{f>f(\Theta)+\varepsilon\}}fd\mu$
$\geq\mu N_{\varepsilon}\cdot f(\Theta)+\mu N_{\varepsilon}^{c}\cdot(f(\Theta)+\varepsilon)$
$=\mu N_{\varepsilon}f(\Theta)+(1-\mu N_{\varepsilon})(f(\Theta)+\varepsilon)$
@@
Therefore, we arrive at the natural conclusion that a measure minimising the expected loss, $P_{n}^{\star}$, is a Dirac delta measure at minimisers of the loss function.
@@
Make Assumptions 1, 2 and 5, and let $\mathcal{Q}=\mathcal{P}(\Theta)$. Then $(P^{\star}_{n})_{n}$ will be a sequence of Dirac measures placing full measure on the set of points minimising the corresponding loss. As a consequence we have that $J_{L_{n}}(P_{n}^{\star})=L_{n}(\Theta)$.
@@
By Lemma 3 we have that the measure minimising $J_{L_{n}}(\cdot)$ will have full measure on the set $N^{n}_{0}:=\{\theta\in\Theta:L(X_{1}^{n},\,\theta)=L_{n}(\Theta)\}$. Since any such measure achieves the lower bound on the functional $J_{L_{n}}(\cdot)$ we may pick any $\theta_{n}^{\star}\in N^{n}_{0}$ and let $P^{\star}_{n}:=\delta_{\theta^{\star}_{n}}$. Therefore, $\delta_{\theta^{\star}_{n}}N_{0}^{n}=1$ as required. ∎
As a result we may characterise convergence of the expected loss minimising measures by the sets $(N_{0}^{n})_{n}$, however noting that the measures are not necessarily unique, even when the space we are minimising over is convex. We may now make the following observation.
@@
Make Assumptions 1, 2, 5 and 7 and suppose $\mathcal{Q}=\mathcal{P}(\Theta)$. Let $A\in\mathcal{T}$ be any set for which $L(A)>L(\Theta)$. Suppose further that for any such $A$ it almost surely holds that
@@
For any such set $A$, there exists an $\varepsilon>0$ such that the sets $N_{\varepsilon}$ and $A$ are disjoint. In particular, by Lemma 3 $\forall n\in\mathbb{N}$ $P_{n}^{\star}N_{\varepsilon}^{n}=1$ $\forall\varepsilon>0$. Therefore, it remains to be shown that eventually almost surely $\exists\varepsilon>0$ such that $A\cap N_{\varepsilon}^{n}=\emptyset$ for all $n$ sufficiently large which will imply the claim.
To see this, by the assumption it is almost sure that
$\liminf_{n}L_{n}(A)>L(\Theta)$
which implies we may pick $\varepsilon=1/2(\liminf_{n}L_{n}(A)-L(\Theta))>0$. Further, as $L_{n}\rightarrow L$ as $n\rightarrow\infty$ pointwise, and $\limsup_{n}L_{n}(\Theta)\leq L(\Theta)$ ($P_{0}$–a.s.) by Assumption 7, so for all $n$ sufficiently large it holds almost surely that $L_{n}(A)-L_{n}(\Theta)>0$. And hence, eventually $P_{0}$–a.s. $A$ and $N_{\varepsilon}^{n}$, with $\varepsilon$ as above, will be disjoint. ∎
5.4.2 A Concentration Result of the GVI Posterior
It is now clear that $\forall\varepsilon>0$ the sequence of loss minimisers $(P_{n}^{\star})_{n}$ concentrates on a set $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}$. Therefore, it seems natural in view of Theorem 6 that the GVI posterior must also concentrate on this set. By Theorem 2 these posteriors exist for all $n$ so we may reason about the behaviour of $(Q_{n})_{n}$. We may explore this rigorously as follows.
@@ 5.4.2 A Concentration Result of the GVI Posterior
Make Assumptions 1, 5, 3, 2, 4, 6 and 7, and suppose that $\mathcal{Q}=\mathcal{P}(\Theta)$. Recall the definition of the set $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}$. Suppose that $\forall\varepsilon>0$ the losses $P_{0}$–a.s. satisfy
@@
We know that $P_{n}^{\star}N_{\varepsilon}\rightarrow 1$ as $n\rightarrow\infty$ $P_{0}$–a.s. $\forall\varepsilon>0$ by Lemma 5. By Theorem 1 we have that $\forall n$ the GVI posteriors $Q_{n}\in\mathcal{R}{{}_{n}^{\star}}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)-J_{L_{n}}(P^{\star}_{n})\leq\frac{M}{\beta n}\}$.
@@
By Theorem 1 we may write for all $n\in\mathbb{N}$
@@
$=Q_{n}N_{\varepsilon}\left(L_{n}(N_{\varepsilon})-L_{n}(\Theta)\right)+Q_{n}N_{\varepsilon}^{c}\left(L_{n}(N_{\varepsilon}^{c})-L_{n}(\Theta)\right)\geq 0\,.$ (12)
@@
We may compare this result with Miller [28] who draws the same conclusion for the Generalised Bayesian posteriors, i.e. GVI with $D=D_{KL}$ and $Q=\mathcal{P}(\Theta)$, with the caveat that there is the requirement on the prior, namely that $\Pi N_{\varepsilon}>0$ which we do not have here. Hence, we may concentrate on areas of the space $\Theta$ which do not have support under the prior, and on which the prior is badly behaved.
5.4.3 Establishing the Concentration of Generalised Variational Inference Posteriors
We may now prove Theorem 6 using these previous lemmas where we first make the following observation. When comparing Lemma 6 with Lemma 5, we note that $A=N_{\varepsilon}^{c}$ is such a set for which $L(A)>L(\Theta)$, and for all such sets $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ there is some $\varepsilon>0$ small enough such that $A\subset N_{\varepsilon}^{c}$.
@@ 5.4.3 Establishing the Concentration of Generalised Variational Inference Posteriors
Fix $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$, then we may pick an $\varepsilon>0$ small enough such that $L(A)>L(\Theta)+\varepsilon$. In particular we may pick $\varepsilon=1/2(L(A)-L(\Theta))>0$, then $A\subset N_{\varepsilon}^{c}$.
@@
Remark.
We have required that $\mathcal{Q}$ allows for Dirac–delta measures here, which we may satisfy by picking $\mathcal{Q}$ the set of Gaussian measures and define the Dirac Delta measures as a Gaussian with 0 covariance; see also Appendix A.
5.5 Rate of Convergence
Recall $N_{\varepsilon}:=\{\theta:L(\theta)\leq L(\Theta)+\varepsilon\}$ for which we know by Theorem 6 that $Q_{n}N_{\varepsilon}^{c}\rightarrow 0$. Now we instead pick a sequence $\varepsilon_{n}\rightarrow 0$ that is non–increasing, and positive, and we aim to show that if $\varepsilon_{n}$ decays sufficiently slowly, then we still have that $Q_{n}N_{\varepsilon_{n}}^{c}\rightarrow 0$. As $N_{\varepsilon_{n}}^{c}\rightarrow\Theta$, hence it will fail that $\liminf_{n}L_{n}(N_{\varepsilon_{n}})>L(\Theta)$. The preceding theory suggests that this should hold if $n\varepsilon_{n}\rightarrow\infty$. In fact, we may now show this under the assumptions of Theorem 7.

### v2

@@ 6.2. Characterising the GVI Posterior through the Empirical Loss Minimiser
Therefore, all probability measures $Q^{\prime}\in\mathcal{Q}$ that satisfy the above bound are potential minimisers of $T_{n}(\cdot)$. This is precisely the set $\mathcal{R}^{\star}_{n}$. Furthermore, by Definition 2 we know that $\forall Q^{\prime}\in\mathcal{Q}$ we have $\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]\geq C$, so the LHS is non–negative, and since $P_{n}^{\star}$ is in this set, it is non–empty.
@@
For the moreover part, we additionally suppose that $\mathcal{Q}$ is weak closed and that the loss is lower semi–continuous and coercive. It immediately follows that the functional $\nu\mapsto J_{L_{n}}(\nu)$ is lower semi–continuous, note that this follows by the functional being over probability measures and the loss being lower bounded and lower semi–continuous [ambrosio2008]. We then know by Lemma 1 that the sublevel sets of the functional $J_{L_{n}}(\cdot)$ are tight. Hence, by the lower semi–continuity of the functional, $\mathcal{Q}$ being weak closed, and Prokhorov’s theorem [billingsley1999], the set $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq k\}$ is weak compact.
@@
Therefore, we may apply this to $\mathcal{R}_{n}^{\star}$ by picking $k=M(n\beta)^{-1}+\inf_{\nu\in\mathcal{Q}}J_{L_{n}}(\nu)$, which exists and is finite by assumption, establishing that $\mathcal{R}_{n}^{\star}$ is weak compact as desired. ∎
6.3. Asymptotics and Frequentist Consistency
@@
Suppose that $f:\Theta\rightarrow\mathbb{R}$ is a $\mathcal{T}$–measurable, lower–bounded, lower semi–continuous, and coercive function. Any minimiser of the functional $J_{f}(\cdot)$ within $\mathcal{P}(\Theta)$ must put full measure on any measurable set containing the minimisers of the function $f(\theta)$.
@@
As the function is lower bounded, lower semi–continuous and coercive, there exists some minimiser away from the boundary of $\Theta$, specifically it achieves its minimum on some compact subset of $\Theta$ away from the boundary, so $N_{0}:=\{\theta\in\Theta:f(\theta)=f(\Theta)\}$ is well defined and non–empty. Note that this is measurable as we may define it as a countable intersection of measurable sets.
Then for any measure $\nu\in\mathcal{P}(\Theta)$ for which $\nu N_{0}=1$ the lower bound above is achieved. Hence, $\forall\varepsilon>0$ and $\forall\mu\in\mathcal{P}(\Theta)$ for which $\mu N_{\varepsilon}<1$, it follows that $J_{f}(\nu)<J_{f}(\mu)$. We may see this as
$J_{f}(\mu)=\int_{N_{\varepsilon}}fd\mu+\int_{N_{\varepsilon}^{c}}fd\mu$ $\geq\mu N_{\varepsilon}\cdot f(\Theta)+\mu N_{\varepsilon}^{c}\cdot(f(\Theta)+\varepsilon)$
@@
Make Assumptions 1, 2 and 3, and let $\mathcal{Q}=\mathcal{P}(\Theta)$. Then $(P^{\star}_{n})_{n}$ will be a sequence of Dirac measures placing full measure on the set of points minimising the corresponding loss. As a consequence we have that $J_{L_{n}}(P_{n}^{\star})=L_{n}(\Theta)$.
@@
By Lemma 3 we have that the measure minimising $J_{L_{n}}(\cdot)$ will have full measure on the set $N^{n}_{0}:=\{\theta\in\Theta:L(X_{1}^{n},\,\theta)=L_{n}(\Theta)\}$. Since any such measure achieves the lower bound on the functional $J_{L_{n}}(\cdot)$ we may pick any $\theta_{n}^{\star}\in N^{n}_{0}$ and let $P^{\star}_{n}:=\delta_{\theta^{\star}_{n}}$. Then $\delta_{\theta^{\star}_{n}}N_{0}^{n}=1$ as required. ∎
@@
Make Assumptions 1, 2, 3 and 6 and suppose $\mathcal{Q}=\mathcal{P}(\Theta)$. Let $A\in\mathcal{T}$ be any set for which $L(A)>L(\Theta)$. Suppose further that for any such $A$ it almost surely holds that
@@
For any such set $A$, there exists an $\varepsilon>0$ such that the sets $N_{\varepsilon}$ and $A$ are disjoint. In particular, by Lemma 3 $\forall n\in\mathbb{N}$ $P_{n}^{\star}N_{\varepsilon}^{n}=1$ $\forall\varepsilon>0$. Therefore, it remains to be shown that eventually almost surely $\exists\varepsilon>0$ such that $A\cap N_{\varepsilon}^{n}=\emptyset$ for all $n$ sufficiently large which will imply the claim. To see this, by the assumption it is almost sure that $\liminf_{n}L_{n}(A)>L(\Theta)$ which implies we may pick $\varepsilon=1/2(\liminf_{n}L_{n}(A)-L(\Theta))>0$. Further, $L_{n}\rightarrow L$ as $n\rightarrow\infty$ pointwise, and $\limsup_{n}L_{n}(\Theta)\leq L(\Theta)$ ($P_{0}$–a.s.) by Assumption 6. Hence for all $n$ sufficiently large it holds almost surely that $L_{n}(A)-L_{n}(\Theta)>0$. And hence, eventually $P_{0}$–a.s. $A$ and $N_{\varepsilon}^{n}$, with $\varepsilon$ as above, will be disjoint. ∎
@@
Make Assumptions 1, 3, 5, 2, 4 and 6, and suppose that $\mathcal{Q}=\mathcal{P}(\Theta)$. Recall the definition of the set $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}$. Suppose that $\forall\varepsilon>0$ the losses $P_{0}$–a.s. satisfy
@@
We know that $P_{n}^{\star}N_{\varepsilon}\rightarrow 1$ as $n\rightarrow\infty$ $P_{0}$–a.s. $\forall\varepsilon>0$ by Lemma 5. By Proposition 5 we have that $\forall n$ the GVI posteriors $Q_{n}\in\mathcal{R}{{}_{n}^{\star}}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)-J_{L_{n}}(P^{\star}_{n})\leq\frac{M}{\beta n}\}$.
@@
By Proposition 5 we may write for all $n\in\mathbb{N}$
@@
(9) $=Q_{n}N_{\varepsilon}\left(L_{n}(N_{\varepsilon})-L_{n}(\Theta)\right)+Q_{n}N_{\varepsilon}^{c}\left(L_{n}(N_{\varepsilon}^{c})-L_{n}(\Theta)\right)\geq 0\,.$
@@
Fix $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$, and pick $\varepsilon=1/2(L(A)-L(\Theta))>0$ for which it holds that $A\subset N_{\varepsilon}^{c}$.
@@
6.4. Consistency under Arbitrary Divergences
Proof of Theorem 7.
By Theorem 1 $Q_{n}$ exists $\forall n$. Then, for any such sequence $(\nu_{n})\subset\mathcal{Q}$ as specified in the assumption it follows that $P_{0}$ almost surely
$n^{-1}T_{n}(\nu_{n})-L_{n}(\Theta)=J_{L_{n}}(\nu_{n})-L_{n}(\Theta)+(n\beta)^{-1}D(\nu_{n}:\Pi)\rightarrow 0\,.$
$\frac{1}{n}T_{n}(\nu_{n})\geq\frac{1}{n}T_{n}(Q_{n})\geq J_{L_{n}}(Q_{n})$ $=\int_{A}L_{n}(\theta)Q_{n}(d\theta)+\int_{A^{c}}L_{n}(\theta)Q_{n}(d\theta)$
$\geq Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(A^{c})\geq L_{n}(\Theta)\,.$
It follows that $n^{-1}T_{n}(Q_{n})-L_{n}(\Theta)\rightarrow 0$. This implies
$J_{L_{n}}(Q_{n})-L_{n}(\Theta)\rightarrow 0\,,\quad\mathrm{and}\quad(n\beta)^{-1}D(Q_{n}:\Pi)\rightarrow 0$
as both quantities are non–negative. Therefore, $J_{L_{n}}(Q_{n})-L_{n}(\Theta)\geq Q_{n}A(L_{n}(A)-L_{n}(\Theta))\geq 0$. But as
$\liminf_{n\rightarrow\infty}L_{n}(A)>L(\Theta)\,,\quad\mathrm{and}\quad\limsup_{n\rightarrow\infty}L_{n}(\Theta)<L(\Theta)$
$\implies\liminf_{n\rightarrow\infty}(L_{n}(A)-L_{n}(\Theta))\geq\liminf_{n\rightarrow\infty}L_{n}(A)-\limsup_{n\rightarrow\infty}L_{n}(\Theta)>0\,$,
and the result follows. ∎
6.5. Generalisation Performance
Proof of Theorem 6.
The proof follows exactly the same steps as Theorem 5 in [shalizi2009] so we do not provide it here. ∎
6.6. Rate of Convergence